# Spectral selectors and contact orderability

### v1

@@ 4.5. Equivalence of the Colin-Sandon oscillation metric and the Fraser-Polterovich-Rosen metric
Let us study the natural generalization of the norm introduced by Fraser-Polterovich-Rosen [35] in the context of Legendrian isotopy classes. Let $(M,\xi)$ be a contact manifold endowed with a contact form $\alpha$ the Reeb flow of which is $1$-periodic. If $\widetilde{\mathcal{G}}$ is orderable, the Fraser-Polterovich-Rosen norm of $\varphi\in\widetilde{\mathcal{G}}$ is defined as
@@
In the original setting of $\widetilde{\mathcal{G}}$, the domination of $\mathrm{d}_{\mathrm{FPR}}$ over the oscillation metric of Colin-Sandon has been remarked by Fraser-Polterovich-Rosen [35, Remark 3.8]. The close link between $\mathrm{d}_{\mathrm{FPR}}$ and the spectral invariant ±$\ell^{\alpha}_{\pm}$ allows us to prove even more in the Legendrian case.
@@
Let $(M,\xi)$ be a contact manifold endowed with a contact form $\alpha$ the Reeb flow of which is $1$-periodic and let us assume that ~$\widetilde{\mathcal{L}}(\Lambda_{*})$ is orderable for some closed Legendrian submanifold $\Lambda_{*}\subset M$. Then the associated Colin-Sandon oscillation metric and Fraser-Polterovich-Rosen metric satisfy
@@
The domination of $\mathrm{d}_{\mathrm{CS,osc}}$ by $\mathrm{d}_{\mathrm{FPR}}$ had already been remarked by Fraser-Polterovich-Rosen [35, Remark 3.8]. It is a consequence of the fact that $\mathrm{d}_{\mathrm{CS,osc}}$ induces a partially ordered metric space on $(\widetilde{\mathcal{L}},\preceq)$ (see [26, Proposition 3.4]). Let us fix $A:=\mathrm{d}_{\mathrm{CS,osc}}(\Lambda_{0},\phi_{1}^{\alpha}\Lambda_{0})$ for some $\Lambda_{0}\in\widetilde{\mathcal{L}}$ and let us show that
@@
where we have used the $\widetilde{\mathcal{G}}$-invariance of $\mathrm{d}_{\mathrm{CS,osc}}$ and the fact that $\phi_{k}^{\alpha}$ commutes with $g^{\pm 1}$. The inequality (19) now follows from the triangular inequality associated with the invariance of the distance under the action of the Reeb flow. Now, let $\Lambda,\Lambda^{\prime}\in\widetilde{\mathcal{L}}$ and let $k:=\mathrm{d}_{\mathrm{FPR}}(\Lambda,\Lambda^{\prime})$. By definition, $\phi_{-k}^{\alpha}\Lambda^{\prime}\preceq\Lambda\preceq\phi_{k}^{\alpha}\Lambda^{\prime}$ so $\mathrm{d}_{\mathrm{CS,osc}}(\Lambda,\phi_{-k}^{\alpha}\Lambda^{\prime})\leq\mathrm{d}_{\mathrm{CS,osc}}(\phi_{-k}^{\alpha}\Lambda^{\prime},\phi_{k}^{\alpha}\Lambda^{\prime})$ by compatibility of $\mathrm{d}_{\mathrm{CS,osc}}$ with the partial order $\preceq$. Therefore,
@@
and make the necessary changes in the definition of $\mathrm{d}_{\mathrm{spec}}^{\alpha}$ as well as $\mathrm{d}_{\mathrm{FPR}}$. As $(\phi_{t})$ is not an autonomous flow, the triangular inequality is not accessible anymore but one still has
@@ 5. Lorentzian geometry and time functions
When $(M,\ker\alpha)$ is a closed cooriented contact manifold, the relations $\preceq$ and on $\mathcal{G}$ (resp. $\widetilde{\mathcal{G}}$) have the property to come from a closed proper cone structure, i.e. a distribution of closed sharp convex cones with non empty interiors [34, 55, 40]. Indeed, the Lie algebra $\mathfrak{g}$ of the infinite dimensional Lie group $\mathcal{G}$ is the space of contact vector fields. The subset $\mathfrak{g}_{\geq 0}$ of contact vector fields $X\in\mathfrak{g}\setminus\{0\}$ satisfying $\alpha(X)\geq 0$ is a closed proper cone whose interior $\mathfrak{g}_{>0}$ consists of vector fields for which the previous inequality is open. The closed proper cone structure used to define $\preceq$ and is then given by right translating these cones of the Lie algebra to the whole tangent space. Note that left translations would give rise to the same closed proper cone structure since $\mathfrak{g}_{\geq 0}$ is invariant under the adjoint action of $\mathcal{G}$ (which is the push forward). This point of view recently led Abbondandolo-Benedetti-Polterovich [1] and Hedicke [41] to introduce objects of Lorentzian geometry to study $(\mathcal{G},\preceq,\cll)$ such as Lorentz-Finsler structure or Lorentzian distances. An open question in [1] was about the existence or not of a time function on $(\widetilde{\mathcal{G}}(\mathbb{R}\mathrm{P}^{2n-1}),\preceq)$, i.e. a function $\tau$ such that $x\preceq y$ and $x\neq y$ implies $\tau(x)<\tau(y)$. In this section, we give a positive answer to this question generalized to all orderable $\mathcal{G},\widetilde{\mathcal{G}},\mathcal{L}$ and $\widetilde{\mathcal{L}}$. We show moreover that time functions cannot be invariant.
@@
Let us fix a contact form $\alpha$ supporting $\xi$. According to Lemma 2.5, $\widetilde{\mathcal{G}}$ is separable (and so is $\mathcal{G}$). Then, let us fix $(\psi_{n})_{n\geq 1}$ a dense sequence on $\mathcal{G}$ (resp. $\widetilde{\mathcal{G}}$) and consider
@@
$\tau^{\alpha}$ is a well-defined $\mathrm{d}_{\mathrm{spec}}^{\alpha}$-$1$-Lipschitz (so $\mathrm{d}_{\mathrm{SH}}^{\alpha}$-$1$-Lipschitz and $C^{1}$-continuous) map $\mathcal{G}\to\mathbb{R}$ (resp. $\widetilde{\mathcal{G}}\to\mathbb{R}$) satisfying
@@
In order to simplify the notation let $c:=c_{+}^{\alpha}$. According to the triangular inequality,
$c(\varphi)-c(\psi^{-1})\leq c(\varphi\psi)\leq c(\varphi)+c(\psi),$
where the left-hand side inequality comes from the decomposition $\varphi=(\varphi\psi)\psi^{-1}$. Therefore,
$\big{|}|c(\varphi\psi)|-|c(\varphi)|\big{|}\leq\max\big{(}|c(\psi)|,|c(\psi^{-1})|\big{)}.$
Applying this inequality with $\psi=\psi_{n}$, one gets that the sum defining $\tau^{\alpha}$ is absolutely convergent. Since $c$ is $\mathrm{d}_{\mathrm{spec}}^{\alpha}$-1-Lipschitz (Equation (15)), and $\mathrm{d}_{\mathrm{spec}}^{\alpha}$ is right-invariant, one gets
@@
One defines time functions on $\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$) similarly. Keeping the previously fixed dense sequence $(\psi_{n})$ of $\mathcal{G}$ (resp. $\widetilde{\mathcal{G}}$) and fixing some $\Lambda_{0}\in\mathcal{L}$ (resp. in $\widetilde{\mathcal{L}})$, let us consider
$\tau^{\alpha}_{\Lambda_{0}}(\Lambda):=a\sum_{n}\frac{\ell^{\alpha}_{+}(\Lambda,\psi_{n}\Lambda_{0})}{2^{n}\max(1,|c_{\pm}^{\alpha}(\psi_{n})|)}+b,\quad\forall\Lambda\in\mathcal{L}\text{ (resp. $\widetilde{\mathcal{L}}$)},$
where $a\in\mathbb{R}$ is defined as above and $b\in\mathbb{R}$ is such that $\tau^{\alpha}_{\Lambda_{0}}(\Lambda_{0})=0$. One proves similarly the Legendrian counterpart of Theorem 5.2
@@
$\tau^{\alpha}_{\Lambda_{0}}$ is a well-defined $\mathrm{d}_{\mathrm{spec}}^{\alpha}$-$1$-Lipschitz (so $\mathrm{d}_{\mathrm{SCH}}^{\alpha}$-$1$-Lipschitz and $C^{1}$-continuous) map $\mathcal{L}\to\mathbb{R}$ (resp. $-~\widetilde{\mathcal{L}}\to\mathbb{R}$) satisfying
@@
    $\tau^{\alpha}_{\Lambda_{0}}(\Lambda_{0})=0$,
@@
    $\tau^{\alpha}_{\Lambda_{0}}(\phi_{t}^{\alpha}\Lambda)=\tau^{\alpha}_{\Lambda_{0}}(\Lambda)+t$ for all $t\in\mathbb{R}$ and $\Lambda\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$),
@@
    $\Lambda\preceq\Lambda^{\prime}$ with $\Lambda\neq\Lambda^{\prime}$ implies $\tau^{\alpha}_{\Lambda_{0}}(\Lambda)<\tau^{\alpha}_{\Lambda_{0}}(\Lambda^{\prime})$.
@@
We will prove that there exist $g,\varphi\in\mathcal{G}$ (resp. $\widetilde{\mathcal{G}}$) such that (20) is satisfied. Let us first remark that it is enough to prove it in the standard contact vector space $(\mathbb{R}^{2n+1},\xi_{\mathrm{st}})$ (we recall that in an open manifold, $\mathcal{G}$ stands for the set of time-one maps of compactly supported contact flows). Indeed, there would exists contact isotopies $(g_{t})$, $(\varphi_{t})$ and $(h_{t})$ supported in some open ball $B\subset\mathbb{R}^{2n+1}$ such that $g_{0}=\mathrm{id}=\varphi_{0}$, $(h_{t})$ is non-negative with $h_{0}=\varphi_{1}$ and $h_{1}=g_{1}^{-1}\varphi_{1}g_{1}$, and $\varphi_{1}\neq g_{1}^{-1}\varphi_{1}g_{1}$. Now, given any contact $(2n+1)$-manifold $(M,\xi)$, there exists a contact embedding of $B$ inside $(M,\xi)$ by the Darboux neighborhood theorem. Since the contact isotopies $(g_{t})$, $(\varphi_{t})$ and $(h_{t})$ are compactly supported in $B$, they naturally extend by the identity to contact isotopies of $(M,\xi)$. Seen in $(M,\xi)$, $(h_{t})$ is still non-negative so that $\varphi:=\varphi_{1}$ and $g:=g_{1}$ still satisfy (20) as needed to conclude.
Let us now prove the existence of such $g,\varphi\in\mathcal{G}$ (resp. $\widetilde{\mathcal{G}}$) for $(\mathbb{R}^{2n+1},\xi_{\mathrm{st}})$ to finish the proof. The standard contact form $\alpha_{\mathrm{st}}$ is $\mathrm{d}z-\sum_{i}y_{i}\mathrm{d}x_{i}$ where $(x,y,z)$ denotes the usual coordinate functions on $\mathbb{R}^{2n+1}$. Let $\rho:[0,+\infty)\to[0,+\infty)$ be a smooth non-increasing function supported in $[0,1/4]$ such that $\rho(0)=1$. Let $H:\mathbb{R}^{2n+1}\to\mathbb{R}$ be the Hamiltonian map defined by $H(p):=\rho(|p|^{2})$ for all $p\in\mathbb{R}^{2n+1}$ and where ||$\cdot$|| ⋅ | denotes the usual Euclidean norm. We denote by $(\varphi_{t})$ the contact flow generated by $H$. Finally for all $a\in\mathbb{R}$ we denote by $\Phi_{a}$ the non compactly supported contactomorphism of $\mathbb{R}^{2n+1}$ defined as $(x,y,z)\mapsto(e^{a}x,e^{a}y,e^{2a}z)$.
Let $a<0$ be sufficiently close to 00 so that $\Phi_{a}^{-1}(B_{0}(1/2))$ is strictly included in $B_{0}(1)$, where $B_{0}(r)$ denotes the open ball centered at 00 of radius $r>0$. Then the support of $\Phi_{a}^{-1}\circ\varphi_{t}\circ\Phi_{a}$, that is $\Phi_{a}^{-1}(\operatorname{supp}\varphi_{t})$, is contained in $B_{0}(1)$ and contains strictly $\operatorname{supp}\varphi_{t}$ for all $t\in[0,1]$. Moreover since the compactly supported Hamiltonian function

### v2

@@ 4.5. Equivalence of the Colin-Sandon oscillation metric and the Fraser-Polterovich-Rosen metric
Let us study the natural generalization of the norm introduced by Fraser-Polterovich-Rosen [36] in the context of Legendrian isotopy classes. Let $(M,\xi)$ be a contact manifold endowed with a contact form $\alpha$ the Reeb flow of which is 111-periodic. If $\widetilde{\mathcal{G}}$ is orderable, the Fraser-Polterovich-Rosen norm of $\varphi\in\widetilde{\mathcal{G}}$ is defined as
@@
In the original setting of $\widetilde{\mathcal{G}}$, the domination of $\mathrm{d}_{\mathrm{FPR}}$ over the oscillation metric of Colin-Sandon has been remarked by Fraser-Polterovich-Rosen [36, Remark 3.8]. The close link between $\mathrm{d}_{\mathrm{FPR}}$ and the spectral invariant ±$\ell^{\alpha}_{\pm}$ allows us to prove even more in the Legendrian case.
@@
Let $(M,\xi)$ be a contact manifold endowed with a contact form $\alpha$ the Reeb flow of which is 111-periodic and let us assume that ~$\widetilde{\mathcal{L}}(\Lambda_{*})$ is orderable for some closed Legendrian submanifold $\Lambda_{*}\subset M$. Then the associated Colin-Sandon oscillation metric and Fraser-Polterovich-Rosen metric satisfy
@@
The domination of $\mathrm{d}_{\mathrm{CS,osc}}$ by $\mathrm{d}_{\mathrm{FPR}}$ had already been remarked by Fraser-Polterovich-Rosen [36, Remark 3.8]. It is a consequence of the fact that $\mathrm{d}_{\mathrm{CS,osc}}$ induces a partially ordered metric space on $(\widetilde{\mathcal{L}},\preceq)$ (see [27, Proposition 3.4]). Let us fix $A:=\mathrm{d}_{\mathrm{CS,osc}}(\Lambda_{0},\phi_{1}^{\alpha}\Lambda_{0})$ for some $\Lambda_{0}\in\widetilde{\mathcal{L}}$ and let us show that
@@
where we have used the $\widetilde{\mathcal{G}}$-invariance of $\mathrm{d}_{\mathrm{CS,osc}}$ and the fact that $\phi_{k}^{\alpha}$ commutes with $g^{\pm 1}$. The inequality (20) now follows from the triangle inequality associated with the invariance of the distance under the action of the Reeb flow. Now, let $\Lambda,\Lambda^{\prime}\in\widetilde{\mathcal{L}}$ and let $k:=\mathrm{d}_{\mathrm{FPR}}(\Lambda,\Lambda^{\prime})$. By definition, $\phi_{-k}^{\alpha}\Lambda^{\prime}\preceq\Lambda\preceq\phi_{k}^{\alpha}\Lambda^{\prime}$ so $\mathrm{d}_{\mathrm{CS,osc}}(\Lambda,\phi_{-k}^{\alpha}\Lambda^{\prime})\leq\mathrm{d}_{\mathrm{CS,osc}}(\phi_{-k}^{\alpha}\Lambda^{\prime},\phi_{k}^{\alpha}\Lambda^{\prime})$ by compatibility of $\mathrm{d}_{\mathrm{CS,osc}}$ with the partial order $\preceq$. Therefore,
@@
and make the necessary changes in the definition of $\mathrm{d}_{\mathrm{spec}}^{\alpha}$ as well as $\mathrm{d}_{\mathrm{FPR}}$. As $(\phi_{t})$ is not an autonomous flow, the triangle inequality is not accessible anymore but one still has
@@ 5. Lorentzian geometry and time functions
When $(M,\ker\alpha)$ is a closed cooriented contact manifold, the relations $\preceq$ and on $\mathcal{G}$ (resp. $\widetilde{\mathcal{G}}$) have the property to come from a closed proper cone structure, i.e. a distribution of closed sharp convex cones with non empty interiors [35, 56, 41]. Indeed, the Lie algebra $\mathfrak{g}$ of the infinite dimensional Lie group $\mathcal{G}$ is the space of contact vector fields. The subset $\mathfrak{g}_{\geq 0}$ of contact vector fields $X\in\mathfrak{g}\setminus\{0\}$ satisfying $\alpha(X)\geq 0$ is a closed proper cone whose interior $\mathfrak{g}_{>0}$ consists of vector fields for which the previous inequality is open. The closed proper cone structure used to define $\preceq$ and is then given by right translating these cones of the Lie algebra to the whole tangent space. Note that left translations would give rise to the same closed proper cone structure since $\mathfrak{g}_{\geq 0}$ is invariant under the adjoint action of $\mathcal{G}$ (which is the push forward). This point of view recently led Abbondandolo-Benedetti-Polterovich [1] and Hedicke [42] to introduce objects of Lorentzian geometry to study $(\mathcal{G},\preceq,\cll)$ such as Lorentz-Finsler structure, Lorentzian distances or time functions. Recall that a continuous real valued function $\tau$ on a partially ordered topological space $(\mathcal{O},\preceq)$ is a time function if $x\preceq y$ and $x\neq y$ implies that $\tau(x)<\tau(y)$. An open question in [1] was about the existence or not of a time function on $(\widetilde{\mathcal{G}}(\mathbb{R}\mathrm{P}^{2n-1}),\preceq)$. In this section, we give a positive answer to this question generalized to all orderable $\mathcal{G},\widetilde{\mathcal{G}},\mathcal{L}$ and $\widetilde{\mathcal{L}}$. We show moreover that time functions cannot be invariant (see Theorem 5.4 for a more precise statement).
@@
Let us fix a contact form $\alpha$ supporting $\xi$ and suppose that $M$ is closed. According to Lemma 2.8, when $M$ is closed, endowed with the $C^{1}$-topology $\widetilde{\mathcal{G}}$ is separable (and so is $\mathcal{G}$). Then, let us fix $(\psi_{n})_{n\in\mathbb{N}}$ a dense sequence on $\mathcal{G}$ (resp. $\widetilde{\mathcal{G}}$) and consider
@@
$\tau^{\alpha}$ is a well-defined $\mathrm{d}_{\mathrm{spec}}^{\alpha}$-111-Lipschitz (so $\mathrm{d}_{\mathrm{SH}}^{\alpha}$-111-Lipschitz and $C^{1}$-continuous) map $\mathcal{G}\to\mathbb{R}$ (resp. $\widetilde{\mathcal{G}}\to\mathbb{R}$) satisfying
@@
In order to simplify the notation let $c:=c_{+}^{\alpha}$ and let $F_{n}=2^{n}\max(1,|c_{\pm}^{\alpha}(\psi_{n})|)=2^{n}\max(1,|c(\psi_{n})|,|c(\psi_{n}^{-1})|)$ for all $n\in\mathbb{N}$. For any $\varphi\in\mathcal{G}$ (resp. $\widetilde{\mathcal{G}}$), according to the triangle inequality,
$c(\varphi)-c(\psi_{n}^{-1})\leq c(\varphi\psi_{n})\leq c(\varphi)+c(\psi_{n}),$
where the left-hand side inequality comes from the decomposition $\varphi=(\varphi\psi_{n})\psi_{n}^{-1}$. Therefore
$|c(\varphi\psi_{n})|$ $\leq\max\{|c(\varphi)+c(\psi_{n})|,|c(\varphi)-c(\psi_{n}^{-1})|\}$ (21)
$\leq\max\{|c(\varphi)|+|c(\psi_{n})|,|c(\varphi)|+|c(\psi_{n}^{-1})|\}.$
Dividing (21) by $F_{n}$ we get $\left|\frac{c(\varphi\psi_{n})}{F_{n}}\right|\leq\frac{1}{2^{n}}\left(|c(\varphi)|+1\right).$ Therefore the series $\sum\limits_{n\in\mathbb{N}}\frac{c(\varphi\psi_{n})}{F_{n}}$ is absolutely converging and thus $\tau^{\alpha}$ as well.
Since $c$ is $\mathrm{d}_{\mathrm{spec}}^{\alpha}$-1-Lipschitz (Equation (16)), and $\mathrm{d}_{\mathrm{spec}}^{\alpha}$ is right-invariant, one gets
@@
One defines similarly time functions on orderable $\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$) . More precisely, take a dense sequence $(\Lambda_{n})$ of $\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$), whose existence is guaranteed by Lemma 2.7 and fix some $\Lambda_{*}\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$). Consider
$\tau^{\alpha}_{\Lambda_{*}}(\Lambda):=a\sum_{n}\frac{\ell^{\alpha}_{+}(\Lambda,\Lambda_{n})}{2^{n}\max(1,|\ell_{\pm}^{\alpha}(\Lambda_{*},\Lambda_{n})|)}+b,\quad\forall\Lambda\in\mathcal{L}\text{ (resp. $\widetilde{\mathcal{L}}$)},$
where $a=\left[\sum_{n}\frac{1}{2^{n}\max(1,|\ell_{\pm}^{\alpha}(\Lambda_{*},\Lambda_{n})|)}\right]^{-1}$ and $b\in\mathbb{R}$ is such that $\tau^{\alpha}_{\Lambda_{*}}(\Lambda_{*})=0$.
One proves similarly the Legendrian counterpart of Theorem 5.2
@@
$\tau^{\alpha}_{\Lambda_{*}}$ is a well-defined $\mathrm{d}_{\mathrm{spec}}^{\alpha}$-111-Lipschitz (so $\mathrm{d}_{\mathrm{SCH}}^{\alpha}$-111-Lipschitz and $C^{1}$-continuous) map $\mathcal{L}\to\mathbb{R}$ (resp. $\widetilde{\mathcal{L}}\to\mathbb{R}$) satisfying
@@
    $\tau^{\alpha}_{\Lambda_{*}}(\Lambda_{*})=0$,
@@
    $\tau^{\alpha}_{\Lambda_{*}}(\phi_{t}^{\alpha}\Lambda)=\tau^{\alpha}_{\Lambda_{*}}(\Lambda)+t$ for all $t\in\mathbb{R}$ and $\Lambda\in\mathcal{L}$ (resp. $\widetilde{\mathcal{L}}$),
@@
    $\Lambda\preceq\Lambda^{\prime}$ with $\Lambda\neq\Lambda^{\prime}$ implies $\tau^{\alpha}_{\Lambda_{*}}(\Lambda)<\tau^{\alpha}_{\Lambda_{*}}(\Lambda^{\prime})$.
@@
[…]
Let $a<0$ be sufficiently close to 0 so that $\Phi_{a}^{-1}(B_{0}(1/2))$ is strictly included in $B_{0}(1)$, where $B_{0}(r)$ denotes the open ball centered at 0 of radius $r>0$. Then the support of $\Phi_{a}^{-1}\circ\varphi_{t}\circ\Phi_{a}$, that is $\Phi_{a}^{-1}(\operatorname{supp}\varphi_{t})$, is contained in $B_{0}(1)$ and contains strictly $\operatorname{supp}\varphi_{t}$ for all $t\in[0,1]$. Moreover since the compactly supported Hamiltonian function